whats the slope of 2x+2y+1=0
Answer by nyc_function(2741)

The slope is the number next to x after y has been isolated.
2x + 2y + 1 = 0
2x + 2y = -1
2y = -2x - 1
y = (-2x - 1)/2
y = -x - 1/2
The slope is -1.